There are several gas cylinders that are under pressure, which are located outside of the maintenance department. As the safety professional, you have been asked a couple of questions regarding this issue. Please provide the correct responses, and discuss your recommendations for any control measures. The volume of the gas cylinder is 25.7 liters and exerts a pressure of 670 mmHg. If part of the gas is used, the pressure drops to 595 mmHg. What would be the remaining volume of gas? One of the cylinder's content is under pressure at 1900 psi (per the gauge) at 70°F. As the day heats up because of the sun, the temperature increases to 105°F. What is the pressure at 105°F?
Answer #1

1. Pressure * Volume = constant

If V is the volume of remaining gas, the euation will be like

P1 x V1 = P2 x V2

670 x25.7 = 595 x V

V = 670x25.7 /595 = 28.93 litres

2.

P / T = constant

P1 /T1 = P2 / T2

changing Temperatures to the units

T1 = 700F = (70-32) x5/9+ 273 = 294.11

T2 = 1050F = ( 105-32)x5/9 + 273 = 313.55

P2 = P1 x T2 / T1 = 1900 x313.55 /294.11 = 2025.62 psi